OpenStudy (anonymous):

write the first 5 terms of the sequence? A sub 1= 10, A sub (n -1) -5

OpenStudy (anonymous):

I am really confused

OpenStudy (anonymous):

did you type everything correctly?

OpenStudy (anonymous):

I cant type the subscripts so it doesn't look the same exactly

OpenStudy (anonymous):

\[a _{1}^{}= 10, a _{n-1}^{}-5\] that is correct

OpenStudy (anonymous):

a 1 =10,an=a n−1 −5

OpenStudy (ranga):

What a difference that makes!

OpenStudy (anonymous):

ok, that makes sense now.

OpenStudy (anonymous):

okay Sorry lol

OpenStudy (anonymous):

Substitute n = 2, you get a sub 2 = (a sub 1) - 5 = 10 - 5 = 5 Now, lets put in n = 3, so a sub 3 = (a sub 2) - 5 = 5-5 = 0 Now, lets out in n = 4, so a sub 4 = (a sub 3) - 5 = 0 - 5 = -5 Finally, n = 5, so a sub 5 = (a sub 4) - 5 = -5 - 5 = -10 So 5th term, a sub 5 = -10

OpenStudy (anonymous):

okay thanks :) but you just sub in the a1 and so forth

OpenStudy (anonymous):

Yes, you put in a sub 1 to find a sub 2. Then put in a sub 2 to get a sub 3, and so on. As I showed you step-by-step.